# Agentic Retrieval-Augmented Generation: A Survey on Agentic RAG

### 04 Feb 2025 v3

@@ Abstract
Large Language Models (LLMs) have revolutionized artificial intelligence (AI) by enabling human-like text generation and natural language understanding. However, their reliance on static training data limits their ability to respond to dynamic, real-time queries, resulting in outdated or inaccurate outputs. Retrieval-Augmented Generation (RAG) has emerged as a solution, enhancing LLMs by integrating real-time data retrieval to provide contextually relevant and up-to-date responses. Despite its promise, traditional RAG systems are constrained by static workflows and lack the adaptability required for multi-step reasoning and complex task management.
Agentic Retrieval-Augmented Generation (Agentic RAG) transcends these limitations by embedding autonomous AI agents into the RAG pipeline. These agents leverage agentic design patterns reflection, planning, tool use, and multi-agent collaboration to dynamically manage retrieval strategies, iteratively refine contextual understanding, and adapt workflows through clearly defined operational structures ranging from sequential steps to adaptive collaboration. This integration enables Agentic RAG systems to deliver unparalleled flexibility, scalability, and context-awareness across diverse applications.
This survey provides a comprehensive exploration of Agentic RAG, beginning with its foundational principles and the evolution of RAG paradigms. It presents a detailed taxonomy of Agentic RAG architectures, highlights key applications in industries such as healthcare, finance, and education, and examines practical implementation strategies. Additionally, it addresses challenges in scaling these systems, ensuring ethical decision-making, and optimizing performance for real-world applications, while providing detailed insights into frameworks and tools for implementing Agentic RAG  [^1] GitHub link: https://github.com/asinghcsu/AgenticRAG-Survey. The GitHub link for this survey is available at: https://github.com/asinghcsu/AgenticRAG-Survey.
@@ 1 Introduction
Large Language Models (LLMs) [1, 2] [3], such as OpenAI’s GPT-4, Google’s PaLM, and Meta’s LLaMA, have significantly transformed artificial intelligence (AI) with their ability to generate human-like text and perform complex natural language processing tasks. These models have driven innovation across diverse domains, including conversational agents [4], automated content creation, and real-time translation. Recent advancements have extended their capabilities to multimodal tasks, such as text-to-image and text-to-video generation [5], enabling the creation and editing of videos and images from detailed prompts [6], which broadens the potential applications of generative AI.
Despite these advancements, LLMs face significant limitations due to their reliance on static pre-training data. This reliance often results in outdated information, hallucinated responses [7], and an inability to adapt to dynamic, real-world scenarios. These challenges emphasize the need for systems that can integrate real-time data and dynamically refine responses to maintain contextual relevance and accuracy.
Retrieval-Augmented Generation (RAG) [8, 9] emerged as a promising solution to these challenges. By combining the generative capabilities of LLMs with external retrieval mechanisms [10], RAG systems enhance the relevance and timeliness of responses. These systems retrieve real-time information from sources such as knowledge bases [11], APIs, or the web, effectively bridging the gap between static training data and the demands of dynamic applications. However, traditional RAG workflows remain limited by their linear and static design, which restricts their ability to perform complex multi-step reasoning, integrate deep contextual understanding, and iteratively refine responses.
The evolution of agents [12] has significantly enhanced the capabilities of AI systems. Modern agents, including LLM-powered and mobile agents [13], are intelligent entities capable of perceiving, reasoning, and autonomously executing tasks. These agents leverage agentic patterns, such as reflection [14], planning [15], tool use, and multi-agent collaboration [16], to enhance decision-making and adaptability.
Furthermore, these agents employ agentic workflow patterns [12, 13], such as prompt chaining, routing, parallelization, orchestrator-worker models, and evaluator-optimizer , to structure and optimize task execution. By integrating these patterns, Agentic RAG systems can efficiently manage dynamic workflows and address complex problem-solving scenarios. The convergence of RAG and agentic intelligence has given rise to Agentic Retrieval-Augmented Generation (Agentic RAG) [14], a paradigm that integrates agents into the RAG pipeline. Agentic RAG enables dynamic retrieval strategies, contextual understanding, and iterative refinement [15], allowing for adaptive and efficient information processing. Unlike traditional RAG, Agentic RAG employs autonomous agents to orchestrate retrieval, filter relevant information, and refine responses, excelling in scenarios requiring precision and adaptability. The overview of Agentic RAG is in figure 1.
This survey explores the foundational principles, taxonomy, and applications of Agentic RAG. It provides a comprehensive overview of RAG paradigms, such as Naïve RAG, Modular RAG, and Graph RAG [16], alongside their evolution into Agentic RAG systems. Key contributions include a detailed taxonomy of Agentic RAG frameworks, applications across domains such as healthcare [17, 18], finance, and education [19], and insights into implementation strategies, benchmarks, and ethical considerations.
@@
The structure of this paper is as follows: Section 2 introduces RAG and its evolution, highlighting the limitations of traditional approaches. Section 3 elaborates on the principles of agentic intelligence and agentic patterns. Section 4 elaborates agentic workflow patterns. Section 5 provides a taxonomy of Agentic RAG systems, including single-agent, multi-agent, and graph-based frameworks. Section 6 provides comparative analysis of Agentic RAG frameworks. Section 7 examines applications of Agentic RAG, while Section 8 discusses implementation tools and frameworks. Section 9 focuses on benchmarks and dataset, and Section 10 concludes with future directions for Agentic RAG systems.
@@ 9 Benchmarks and Datasets
Current benchmarks and datasets provide valuable insights into evaluating Retrieval-Augmented Generation (RAG) systems, including those with agentic and graph-based enhancements. While some are explicitly designed for RAG, others are adapted to test retrieval, reasoning, and generation capabilities in diverse scenarios. Datasets are crucial for testing the retrieval, reasoning, and generation components of RAG systems. Table 3 discusses some key datasets based on the dowstream task for RAG Evaluation.
Benchmarks play a critical role in standardizing the evaluation of RAG systems by providing structured tasks and metrics. The following benchmarks are particularly relevant:
@@
  • BEIR (Benchmarking Information Retrieval): A versatile benchmark designed for evaluating embedding models on a variety of information retrieval tasks, encompassing 17 datasets across diverse domains like bioinformatics, finance, and question answering [55].
@@
  • MS MARCO (Microsoft Machine Reading Comprehension): Focused on passage ranking and question answering, this benchmark is widely used for dense retrieval tasks in RAG systems [56].
@@
  • TREC (Text REtrieval Conference, Deep Learning Track): Provides datasets for passage and document retrieval, emphasizing the quality of ranking models in retrieval pipelines [57].
@@
  • MuSiQue (Multihop Sequential Questioning): A benchmark for multihop reasoning across multiple documents, emphasizing the importance of retrieving and synthesizing information from disconnected contexts [58].
  • 2WikiMultihopQA: A dataset designed for multihop QA tasks over two Wikipedia articles, focusing on the ability to connect knowledge across multiple sources [59].
@@
  • AgentG (Agentic RAG for Knowledge Fusion): Tailored for agentic RAG tasks, this benchmark assesses dynamic information synthesis across multiple knowledge bases [8].
@@
  • HotpotQA: A multi-hop QA benchmark requiring retrieval and reasoning over interconnected contexts, ideal for evaluating complex RAG workflows[60].
@@
  • RAGBench: A large-scale, explainable benchmark featuring 100,000 examples across industry domains, with a TRACe evaluation framework for actionable RAG metrics [61].
@@
  • BERGEN (Benchmarking Retrieval-Augmented Generation): A library for systematically benchmarking RAG systems with standardized experiments [62].
@@
  • FlashRAG Toolkit: Implements 12 RAG methods and includes 32 benchmark datasets to support efficient and standardized RAG evaluation [63].
@@
  • GNN-RAG: This benchmark evaluates graph-based RAG systems on tasks like node-level and edge-level predictions, focusing on retrieval quality and reasoning performance in Knowledge Graph Question Answering (KGQA) [64].

### 01 Apr 2026 v4

@@ Abstract
Large Language Models (LLMs) have significantly advanced artificial intelligence by enabling human-like text generation and natural language understanding. However, their reliance on static training data limits their ability to respond to dynamic, real-time queries, resulting in outdated or inaccurate outputs. Retrieval-Augmented Generation (RAG) has emerged as a solution, enhancing LLMs by integrating real-time data retrieval to provide contextually relevant and up-to-date responses. Despite its promise, traditional RAG systems are constrained by static workflows and lack the adaptability required for multi-step reasoning and complex task management. Agentic Retrieval-Augmented Generation (Agentic RAG) transcends these limitations by embedding autonomous AI agents into the RAG pipeline. These agents leverage agentic design patterns reflection, planning, tool use, and multi-agent collaboration to dynamically manage retrieval strategies, iteratively refine contextual understanding, and adapt workflows through clearly defined operational structures ranging from sequential steps to adaptive collaboration. This integration enables Agentic RAG systems to deliver unparalleled flexibility, scalability, and context-awareness across diverse applications. This paper presents an analytical survey of Agentic RAG [^1] GitHub link: https://github.com/asinghcsu/AgenticRAG-Survey systems. It traces the evolution of RAG paradigms, introduces a principled taxonomy of Agentic RAG architectures based on agent cardinality, control structure, autonomy, and knowledge representation, and provides a comparative analysis of design trade-offs across existing frameworks. The survey further examines real-world applications in domains including healthcare, finance, education, and enterprise document processing, and distills practical lessons learned for system designers and practitioners. Finally, it identifies key open research challenges related to evaluation, coordination, memory management, efficiency, and governance, outlining directions for future research.
@@ 1 Introduction
Large Language Models (LLMs) [57, 80, 103], such as GPT-5, PaLM, and LLaMA, have transformed artificial intelligence by enabling human-like text generation and natural language understanding. These models drive innovation in conversational agents [15], content creation, real-time translation, and multimodal generation tasks including text-to-image and text-to-video synthesis [79, 75]. However, a central challenge remains: how LLMs can reliably reason over rapidly changing external knowledge rather than relying solely on static pre-training.
Despite these advances, LLMs remain constrained by static pre-training data, often producing outdated responses, hallucinated content [35], and failing to adapt to evolving real-world information.
Retrieval-Augmented Generation (RAG) [48, 102] addresses these limitations by integrating external knowledge into generation at inference time. By combining LLM capabilities with retrieval mechanisms [39] over vector databases [29], APIs, or web sources, RAG improves factual grounding, relevance, and timeliness. However, traditional RAG pipelines are typically static and linear, limiting complex multi-step reasoning, deep contextual understanding, and iterative response refinement. These limitations motivate autonomous control mechanisms that dynamically decide when to retrieve, how to reformulate queries, and when sufficient evidence has been collected.
In parallel, advances in agentic AI [72] have introduced intelligent agents capable of perception, reasoning, and autonomous task execution [2]. These agents leverage design patterns—reflection, planning, tool use, and multi-agent collaboration—to dynamically decompose tasks and adapt strategies. Such capabilities are directly relevant to RAG, as retrieval and reasoning decisions can themselves be treated as agentic tasks. These patterns are further structured through workflow patterns [2, 4], including prompt chaining, routing, parallelization, orchestrator–worker models, and evaluator–optimizer loops, governing how multiple prompts, models, or agents are coordinated at the system level.
The convergence of RAG and agentic intelligence has given rise to Agentic Retrieval-Augmented Generation (Agentic RAG) [65], which integrates autonomous agents directly into the RAG pipeline to enable dynamic retrieval, iterative context refinement, and adaptive workflow orchestration [34]. This introduces an explicit control layer that guides how the system reasons over external evidence—coordinating retrieval, validating information, invoking tools, and refining responses—making Agentic RAG effective for complex, multi-domain tasks requiring precision and adaptability, as illustrated in Fig. 1.
Despite rapidly growing interest, the Agentic RAG field remains fragmented, with diverse architectures, inconsistent terminology, and limited comparative understanding. Existing work focuses on individual frameworks or application-specific implementations, and no unified taxonomy currently organizes Agentic RAG systems along principled design dimensions.
This paper presents an analytical survey aimed at structuring this emerging field. It reviews the evolution of RAG paradigms, including naïve, modular, and graph-based RAG [62], and their transition toward agentic systems. A principled taxonomy of Agentic RAG architectures is introduced based on agent cardinality, control structure, autonomy, and knowledge representation, followed by a comparative analysis of design trade-offs. The survey examines applications across healthcare [78, 28], finance, and education [76], distills practical lessons for system designers, and identifies open research challenges to guide future work.
@@
The structure of this paper is as follows: Section 2 introduces RAG and its evolution, highlighting the limitations of traditional approaches. Section 3 elaborates on the principles of agentic intelligence and agentic patterns. Section 4 elaborates on agentic workflow patterns. Section 5 provides a taxonomy of Agentic RAG systems, including single-agent, multi-agent, and graph-based frameworks. Section 6 provides a comparative analysis of Agentic RAG frameworks. Section 7 examines applications of Agentic RAG, while Section 8 discusses implementation tools and frameworks. Section 9 provides a comparative analysis of Agentic RAG frameworks. Section 10 discusses the lessons learned and practical guidance. Section 11 focuses on benchmarks and datasets, Section 12 discusses open research issues and future challenges, and Section 13 concludes with future directions for Agentic RAG systems.
@@ 8 Tools and Frameworks for Agentic RAG
9 Comparative Analysis of Agentic RAG Frameworks
Table 2 compares Traditional RAG and Agentic RAG across key dimensions. Traditional RAG offers simplicity for basic tasks, while Agentic RAG introduces enhanced reasoning and scalability through multi-agent collaboration. Understanding their respective strengths and limitations is essential for selecting an architecture that meets specific application requirements.
9.1 Comparative Insights and Design Trade-offs
A taxonomy-driven analysis reveals key trade-offs beyond feature-level distinctions. Single-agent architectures favor simplicity and low latency but struggle with multi-domain reasoning. Multi-agent systems improve scalability through parallelism at the cost of coordination overhead. Hierarchical and corrective architectures enhance reliability through strategic oversight but incur higher latency. Graph-based systems enable structured relational reasoning while introducing dependencies on knowledge quality. Table 3 maps these architectures to taxonomy dimensions and design trade-offs.
Table 3: Taxonomy-Driven Comparison of Agentic RAG Architectures
Dimension Single-Agent RAG Multi-Agent RAG Hierarchical / Corrective RAG Graph-Based / Document-Centric RAG
Agent Cardinality Single Multiple Multiple (tiered) Multiple
Control Structure Centralized Flat Hierarchical / evaluator-based Adaptive / structured
Knowledge Representation Unstructured Unstructured or hybrid Unstructured or hybrid Graph-based or document-centric
Knowledge Dependency Low Moderate Moderate High, sensitive to graph quality
Reasoning Mode Implicit, LLM-driven Distributed across agents Strategic with supervisory control Relational or symbolic + LLM
Reasoning Depth Low to moderate Moderate to high High High, multi-hop reasoning
Retrieval Adaptivity Limited High High High
Scalability Moderate High Moderate Moderate
Latency Low Moderate High Moderate to high
Engineering Complexity Low Moderate High High
Best-Fit Scenarios Simple QA and routing Multi-domain synthesis High-stakes, reliability-critical tasks Knowledge-intensive, relational domains
10 Lessons Learned and Practical Guidance
Based on the analysis of paradigms, taxonomies, and real-world use cases, this section distills practical insights for designing, evaluating, and deploying Agentic RAG systems.
10.1 Agentic RAG Is Not Always the Right Default
Agentic RAG should not be viewed as a universal replacement for traditional RAG. While it offers superior adaptability and multi-step reasoning, it also introduces coordination complexity, latency, and computational cost. For simple fact retrieval or well-scoped queries, modular RAG pipelines provide sufficient performance with lower overhead. Practitioners should adopt agentic designs selectively, guided by task complexity.
10.2 Architectural Choice Strongly Shapes System Behavior
Different architectures lead to fundamentally different behaviors. Single-agent systems favor simplicity; multi-agent systems enable parallelism but require orchestration; hierarchical architectures introduce oversight at the expense of complexity. Architectural decisions implicitly encode assumptions about control, trust, and error tolerance—corrective workflows improve accuracy but add latency, while adaptive architectures trade correctness guarantees for responsiveness.
10.3 Retrieval Quality Remains the Primary Bottleneck
Agentic reasoning cannot compensate for consistently poor retrieval. Failures often originate from inadequate retrieval coverage, poorly constructed indexes, or insufficient integration of structured and unstructured knowledge. Investing in robust retrieval pipelines and high-quality indexing before introducing agentic complexity remains essential.
10.4 Agent Autonomy Requires Explicit Constraints
Unrestricted autonomy risks excessive tool invocation, redundant reasoning loops, or misaligned objectives. Effective systems balance autonomy with bounded planning horizons, predefined tool access policies, and explicit stopping criteria. Constrained autonomy yields more reliable outcomes, especially in production environments.
10.5 Evaluation Must Account for Process, Not Just Outcomes
Existing benchmarks focus on output quality with limited visibility into intermediate decisions. Meaningful evaluation requires process-level metrics capturing reasoning efficiency, tool usage patterns, and adaptation to changing contexts. Without such evaluation, improvements risk being anecdotal rather than systematic.
10.6 Domain Knowledge Significantly Amplifies Agentic Benefits
Agentic RAG achieves its strongest gains in domains with structured knowledge and explicit constraints. Healthcare, finance, and legal analysis particularly benefit from combining retrieval with rule-based reasoning and graph-structured knowledge. Open-domain tasks show more modest gains, underscoring domain modeling as a complementary component of Agentic RAG design.
10.7 Toward Responsible Deployment of Agentic RAG
Agentic RAG introduces challenges in transparency, accountability, and trust. Multi-agent collaboration complicates error attribution, while autonomous tool use raises safety concerns. Responsible deployment requires governance mechanisms, human oversight, and clear operational boundaries. Explainability, traceability, and auditability must be first-class design goals, particularly for high-stakes applications.
@@ 11 Benchmarks and Datasets
Current benchmarks [22] and datasets evaluate RAG systems across retrieval, reasoning, and generation. Table 5 summarizes key datasets by downstream task. The following benchmarks are particularly relevant:
@@
  • BEIR (Benchmarking Information Retrieval): A benchmark for evaluating embedding models across 17 datasets in bioinformatics, finance, and question answering [86].
@@
  • MS MARCO (Microsoft Machine Reading Comprehension): Passage ranking and question answering, widely used for dense retrieval evaluation in RAG systems [3].
@@
  • TREC (Text REtrieval Conference, Deep Learning Track): Provides datasets for passage and document retrieval, emphasizing ranking quality in retrieval pipelines [13].
@@
  • MuSiQue (Multihop Sequential Questioning): A benchmark for multi-document reasoning, emphasizing retrieval and synthesis across contexts [88].
  • 2WikiMultihopQA: Designed for multihop QA over Wikipedia articles with cross-source integration[33].
@@
  • AgentG (Agentic RAG for Knowledge Fusion) [48]: Tailored for agenticRAG tasks, this benchmark assesses dynamic information synthesis across knowledge bases.
    Tailored for Agentic RAG, assessing dynamic information synthesis across knowledge bases.
@@
  • HotpotQA: A multihop QA benchmark requiring retrieval and reasoning over interconnected contexts, suitable for evaluating complex RAG workflows [98].
@@
  • RAGBench: A large-scale, explainable benchmark with 100,000 examples across domains, using the TRACe framework for actionable RAG evaluation [23].
@@
  • BERGEN (Benchmarking Retrieval-Augmented Generation): A library for systematically benchmarking RAG systems through standardized experiments [64].
@@
  • FlashRAG Toolkit: Implements multiple RAG methods and benchmark datasets to support efficient and standardized evaluation [40].
@@
  • GNN-RAG: Evaluates graph-based RAG systems on node- and edge-level prediction tasks, focusing on retrieval quality and reasoning performance in knowledge graph question answering [53].
Table 4: Downstream Tasks and Datasets for RAG Evaluation (Adapted from [24])
Category Task Type Datasets and References
Multi-hop QA HotpotQA [98], 2WikiMultiHopQA [33], MuSiQue [88]
Long-form QA ELI5 [21], NarrativeQA (NQA) [44], ASQA [83], QMSum [104]
Domain-specific QA Qasper [16], COVID-QA [58], CMB/MMCU Medical [93]
Multi-choice QA QuALITY [61], ARC, CommonsenseQA [84]
Graph-based QA Graph QA GraphQA [32]
Event Argument Extraction WikiEvent [49], RAMS [20]
Dialog Open-domain Dialog Wizard of Wikipedia (WoW) [18]
Personalized Dialog KBP [92], DuleMon [96]
Task-oriented Dialog CamRest [94]
Reasoning Commonsense Reasoning HellaSwag [99], CommonsenseQA [84]
Chain-of-Thought Reasoning CoT Collection [43]
Complex Reasoning CSQA [70]
Others Language Understanding MMLU, WikiText-103 [46]
Fact Checking and Verification FEVER [87], PubHealth [45]
Strategy QA StrategyQA [25]
Summarization Text Summarization WikiASP [30], XSum [59]
Long-form Summarization NarrativeQA (NQA) [44], QMSum [104]
Text Generation Biography Biography Dataset
Text Classification Sentiment Analysis SST-2 [81]
General Classification VioLens [71], TREC [13]
Code Search Programming Search CodeSearchNet [37]
Robustness Retrieval Robustness NoMIRACL [85]
Language Modeling Robustness WikiText-103 [54]
Math Math Reasoning GSM8K [10]
Machine Translation Translation Tasks JRC-Acquis [82]
12 Open Research Issues and Future Challenges in Agentic RAG
Despite rapid progress in Agentic RAG, the field remains at an early stage of maturity. While existing systems demonstrate clear advantages over traditional RAG pipelines, they also expose a range of unresolved research challenges that must be addressed to enable robust, scalable, and trustworthy deployment in real-world settings. This section outlines key open research issues and articulates a long-term research vision for Agentic RAG systems.
12.1 Agent Coordination, Control, and Emergent Behavior
As systems scale to multi-agent and hierarchical frameworks, emergent behaviors become difficult to predict. Current orchestration relies on heuristic rules and prompt engineering with limited convergence guarantees. Future research must address coordination under partial observability and conflicting objectives, formally model inter-agent dependencies, prevent cascading failures, and ensure verifiable collaborative outcomes.
12.2 Evaluation Methodologies Beyond Output Quality
Output-level metrics (answer correctness, retrieval accuracy) are insufficient for Agentic RAG, where intermediate decisions shape performance. Standardized benchmarks evaluating reasoning trajectories, planning depth, adaptability, robustness under noisy retrieval, and cost efficiency are needed. Process-aware evaluation assessing how answers are produced is essential for systematic progress.
12.3 Memory Management and Long-Term Adaptation
Long-term memory design remains open challenge. Persistent memory risks knowledge drift and bias reinforcement, while frequent updates can amplify hallucinations. Key questions include balancing persistence with adaptability, selective retention, and reconciling external knowledge with stored agent experiences.
12.4 Computational Cost, Efficiency, and Sustainability
Multi-agent collaboration and iterative retrieval increase latency and resource consumption. Future research must explore cost-aware planning, adaptive inference, and lightweight coordination—dynamically adjusting active agents, selectively invoking tools, and optimizing retrieval depth based on task complexity.
12.5 Safety, Trust, and Governance in Autonomous RAG Systems
Autonomous agents may introduce failure modes including unauthorized actions and biased decisions. In multi-agent settings, tracing responsibility is particularly challenging. Open directions include transparent decision-tracing, human-in-the-loop oversight, and governance frameworks ensuring alignment between agent objectives and human intent in high-stakes applications.
12.6 Generalization Across Domains and Tasks
Most existing Agentic RAG systems are evaluated in domain-specific settings or carefully curated benchmarks. It remains unclear how well these systems generalize across domains with differing data characteristics, regulatory constraints, or reasoning requirements. Designing agents that can adapt their retrieval and reasoning strategies without extensive task-specific tuning is an open problem. Future work should focus on developing domain-agnostic agentic abstractions, transferable planning strategies, and adaptive retrieval policies that generalize beyond narrow application contexts. Achieving robust cross-domain generalization will be essential for realizing the full potential of Agentic RAG as a general-purpose reasoning framework.
13 Conclusion
Agentic RAG represents a significant evolution beyond static RAG pipelines toward autonomous, adaptive reasoning frameworks. This survey examined RAG foundations, agentic intelligence integration, and the architectural patterns defining modern Agentic RAG systems. By organizing approaches through a principled taxonomy based on agent cardinality, control structure, autonomy, and knowledge representation, this work provides a structured view of a fragmented landscape. The analysis demonstrates that no single architecture is universally optimal; effectiveness depends on aligning complexity with task requirements. Agentic RAG offers clear advantages for complex, knowledge-intensive tasks in healthcare, finance, education, and enterprise processing, but these benefits require high-quality retrieval foundations, disciplined autonomy, and careful orchestration.
Substantial open challenges remain in multi-agent coordination, process-aware evaluation, memory management, computational efficiency, and governance. This survey consolidates existing work, introduces a unifying framework, and articulates design trade-offs and research directions to advance the field of Agentic RAG.